### v3

@@ 5.2 The Algorithm
Algorithm 1 The algorithm solving APC from [CLL13].
1:Compute the matrix $M=(I-A)^{-1}$.
2:For each pair of vertices $(s,t)$, return
$\operatorname{rank}M[E_{\text{out}}(s),E_{\text{in}}(t)]$
as the value of $\lambda(s,t)$.
@@ 6 Bounded Connectivity
The first idea is to reduce degrees in $G$ while preserving the values of small connectivities. In Section 6.1, we present a simple transformation (from [AJ23, Section 5]) which decreases the degrees of nodes in $G$ to $k$, while preserving the $\min(k,\lambda(s,t))$ values. Following this modification, we only need to compute ranks of $k\times k$ submatrices in step 2 above.
@@ 6.2 Low-Rank Edge-Adjacency
$\det\leavevmode\nobreak\ (I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C})\right).$ (17)
@@
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\tilde{\xi}(\mathcal{C})\right).$ (18)
@@
Figure 3: Given edge-disjoint paths $P_{1}=\langle e_{1},f_{1}\rangle$ and $P_{2}=\langle e_{2},f_{2}\rangle$ in $G$, the determinant of the matrix $(I-X)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ enumerates this pair via the monomial $\xi(P_{1},P_{2})=x_{e_{1}f_{1}}\cdot x_{e_{2}f_{2}}$. The variables in this monomial provide enough information to uniquely recover $P_{1}$ and $P_{2}$. In contrast, for $k=2$, the determinant of $(I-YZ)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ assigns this pair weight $\tilde{\xi}(P_{1},P_{2})=(y_{e_{1}1}z_{1f_{1}}+y_{e_{1}2}z_{2f_{1}})(y_{e_{2}1}z_{1f_{2}}+y_{e_{2}2}z_{2f_{2}}).$ One of the terms in the expansion of the above product is $y_{e_{1}1}z_{1f_{1}}\cdot y_{e_{2}2}z_{2f_{2}}$. We can read this term as saying “the first path $P_{1}$ traverses $e_{1}$ and $f_{1}$, and the second path $P_{2}$ traverses $e_{2}$ and $f_{2}$.” So this monomial provides enough information to recover the pair of paths $\langle P_{1},P_{2}\rangle$ as well.
@@ 6.3 Random Evaluation
$(I-YZ)^{-1}[S,T]=\frac{(\operatorname{adj}(I-YZ))[S,T]}{\det\leavevmode\nobreak\ (I-YZ)}.$
@@
$\det\leavevmode\nobreak\ (I-YZ)^{-1}[S,T]=\frac{\det\leavevmode\nobreak\ (\operatorname{adj}(I-YZ))[S,T]}{Q^{r}}.$ (22)
@@
$\operatorname{rank}\leavevmode\nobreak\ (\tilde{B}(I-CB)^{-1}\tilde{C})[E_{\text{out}}(s),E_{\text{in}}(t)]=\min(k,\lambda(s,t)).$
@@
$\operatorname{rank}\leavevmode\nobreak\ (\tilde{B}(I-CB)^{-1}\tilde{C})[E_{\text{out}}(s),E_{\text{in}}(t)]=\min(k,\lambda(s,t))$
@@ 6.4 The Algorithm
Algorithm 2 The algorithm solving $k$-APC from [AJ23].
1:Compute the matrix $M=\tilde{B}(I-CB)^{-1}\tilde{C}$.
2:For each pair $(s,t)$ of original vertices, return
@@
In step 1 of Algorithm 2, we compute $\tilde{B}(I-CB)^{-1}\tilde{C}$. The matrix $CB$ has rows and columns indexed by pairs $(v,i)\in V_{\text{new}}\times[k]$. For any $(u,i),(v,j)\in V_{\text{new}}\times[k]$, we have
@@ 7 Conclusion
In this paper, we presented alternate derivations of the $\tilde{O}(m^{\omega})$ time algorithm for APC of [CLL13], and the $\tilde{O}((kn)^{\omega})$ time algorithm for $k$-APC of [AJ23]. Our approach works by testing, via random evaluation, whether certain generating functions enumerating edge-disjoint families of paths are nonzero or not.
We conclude this paper by pointing out some connections between this perspective and other classical arguments in mathematics and computer science, and highlighting the current main open problems concerning the complexity of computing connectivities.
@@
Our proof for Lemma 4.7 obtains a generating function for edge-disjoint walks by pairing up monomials corresponding to intersecting walks, and arguing their contributions cancel. This reasoning is essentially identical to the proof of the classic Lindström-Gessel-Viennot lemma from combinatorics, which is often used in mathematics to enumerate families of disjoint lattice paths. We refer the reader to [AZ18, Chapter 29] for an accessible exposition of this theorem and some of its applications. For additional examples of this technique of pairing up and cancelling extraneous terms in matrix algebra, see [Zei85].
@@
Our exposition of the APC and $k$-APC algorithms works by interpreting certain determinants combinatorially, as generating functions for families of edge-disjoint walks. Previous work has also leveraged combinatorial interpretations of the determinant (as polynomials instead of formal power series however) to construct interesting arithmetic circuits [MV97, Cur22]. In particular, the sign-reversing involution designed in [MV97, Proof of Theorem 1] is very similar to the suffix swapping argument we use in the proof of Lemma 4.7.
The general approach of solving graph problems by testing whether certain enumerating polynomials are nonzero is now a common technique in graph algorithms: we refer the reader to [EKW24] and the citations therein for previous examples of this paradigm in the literature.
@@
In the 4-Clique problem, we are given an undirected graph $G$ on $n$ nodes, and are tasked with determining if $G$ contains four vertices which are mutually adjacent. The best conditional lower bounds for APC come from observing that 4-Clique reduces to APC [AGI+19, Section 4]. Using the current best algorithms for rectangular matrix multiplication [WXXZ23, Table 1], the fastest known algorithm for 4-Clique takes $O(n^{3.251})$ time [EG04, Theorem 1]. Even if we conjecture this runtime time is optimal, the resulting lower bound for APC is a far cry from the $n^{4+o(1)}$ runtime we have for APC in dense graphs. Moreover, if $\omega=2$ this lower bound for APC weakens to $n^{3-o(1)}$.
@@
The fastest known deterministic verifier for APC runs in $O(n^{3.251}+n^{2.5}\sqrt{m})$ time [Tra23]. It would be interesting to improve this runtime to $O(n^{3.251})$ to match the fastest known algorithm (and deterministic verifier runtime) for 4-Clique.
It would also be interesting to obtain faster randomized verifiers for APC. Currently, no randomized verifiers running faster than the deterministic verifier discussed above are known for APC. This is surprising, since the best lower bound for APC comes from the 4-Clique problem, and the 4-Clique problem admits a randomized verifier running in near-optimal $\tilde{O}(n^{2})$ time [ACJ+23, Section 4].
@@
In directed graphs, APVC and $k$-APVC reduce to APC and $k$-APVC respectively, so it might be easier to design faster algorithms for the former problems instead of the latter problems. For example, it is known that $k$-APVC can be solved in $\tilde{O}(k^{2}n^{\omega})$ time [AJ23, Theorem 5], which is faster than the $\tilde{O}((kn)^{\omega})$ runtime for $k$-APC if $\omega>2$.
In undirected graphs, although APC can be solved in $\tilde{O}(n^{2})$ time, APVC is not known to admit a near-quadratic time algorithm. In fact, the best known lower bounds for APC in directed graphs also hold for APVC in undirected graphs [HLSW23].
So again, it might be easier to find faster algorithms for APVC in undirected graphs instead of tackling the general problem of APC in directed graphs. For example, APVC in undirected graphs can be solved in $\tilde{O}(m^{2})$ time [Tra23, Section 3], which is faster than the $\tilde{O}(m^{\omega})$ runtime for APC if $\omega>2$, and admits a deterministic $O(n^{3.251})$ time verifier [Tra23, Lemma 2.2], which is faster than the known $O(n^{3.521}+n^{2.5}\sqrt{m})$ time deterministic verifier for APC.

### v4

@@ 6 Bounded Connectivity
The first idea is to reduce degrees in $G$ while preserving the values of small connectivities. In Section 6.1, we present a simple transformation (from [k-APC, Section 5]) which decreases the degrees of nodes in $G$ to $k$, while preserving the $\min(k,\lambda(s,t))$ values. Following this modification, we only need to compute ranks of $k\times k$ submatrices in step 2 above.
@@ 6.2 Low-Rank Edge-Adjacency
$\det~{}(I-X)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\xi(\mathcal{C})\right).$ (17)
@@
$\det~{}(I-YZ)^{-1}[S,T]=\sum_{\ell=1}^{\infty}\left(\sum_{\mathcal{C}\in\mathcal{D}_{\ell}(S,T)}\tilde{\xi}(\mathcal{C})\right).$ (18)
@@
Figure 3: Given edge-disjoint paths $P_{1}=\langle e_{1},f_{1}\rangle$ and $P_{2}=\langle e_{2},f_{2}\rangle$ in $G$, the determinant of the matrix $(I-X)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ enumerates this pair via the monomial $\xi(P_{1},P_{2})=x_{e_{1}f_{1}}\cdot x_{e_{2}f_{2}}$. The variables in this monomial provide enough information to uniquely recover $P_{1}$ and $P_{2}$. In contrast, for $k=2$, the determinant of $(I-YZ)^{-1}[\left\{e_{1},e_{2}\right\},\left\{f_{1},f_{2}\right\}]$ assigns this pair weight
$\tilde{\xi}(P_{1},P_{2})=(y_{e_{1}1}z_{1f_{1}}+y_{e_{1}2}z_{2f_{1}})(y_{e_{2}1}z_{1f_{2}}+y_{e_{2}2}z_{2f_{2}}).$
One of the terms in the expansion of the above product is $y_{e_{1}1}z_{1f_{1}}\cdot y_{e_{2}2}z_{2f_{2}}$. We can read this term as saying “the first path $P_{1}$ traverses $e_{1}$ and $f_{1}$, and the second path $P_{2}$ traverses $e_{2}$ and $f_{2}$.” So this monomial provides enough information to recover the pair of paths $\langle P_{1},P_{2}\rangle$ as well.
@@ 6.3 Random Evaluation
$(I-YZ)^{-1}[S,T]=\frac{(\operatorname{adj}(I-YZ))[S,T]}{\det~{}(I-YZ)}.$
@@
$\det~{}(I-YZ)^{-1}[S,T]=\frac{\det~{}(\operatorname{adj}(I-YZ))[S,T]}{Q^{r}}.$ (22)
@@
$\operatorname{rank}~{}(\tilde{B}(I-CB)^{-1}\tilde{C})[E_{\text{out}}(s),E_{\text{in}}(t)]=\min(k,\lambda(s,t)).$
@@
$\operatorname{rank}~{}(\tilde{B}(I-CB)^{-1}\tilde{C})[E_{\text{out}}(s),E_{\text{in}}(t)]=\min(k,\lambda(s,t))$
@@ 6.4 The Algorithm
1 Compute the matrix $M=\tilde{B}(I-CB)^{-1}\tilde{C}$.
For each pair $(s,t)$ of original vertices, return
@@
In step 1 of Algorithm 2, we compute $\tilde{B}(I-CB)^{-1}\tilde{C}$. The matrix $CB$ has rows and columns indexed by pairs $(v,i)\in V_{\text{new}}\times[k]$. For any $(u,i)$ and $(v,j)$ in $V_{\text{new}}\times[k]$, we have
@@ 7 Conclusion
In this paper, we presented alternate derivations of the $\tilde{O}(m^{\omega})$ time algorithm for APC of [CheungLauLeung2013], and the $\tilde{O}((kn)^{\omega})$ time algorithm for $k$-APC of [k-APC]. Our approach works by testing, via random evaluation, whether certain generating functions enumerating edge-disjoint families of paths are nonzero or not. We conclude this paper by pointing out some connections between this perspective and other classical arguments in mathematics and computer science, and highlighting the current main open problems concerning the complexity of computing connectivities.
@@
Our proof for Lemma 4.7 obtains a generating function for edge-disjoint walks by pairing up monomials corresponding to intersecting walks, and arguing their contributions cancel. This reasoning is essentially identical to the proof of the classic Lindström-Gessel-Viennot lemma from combinatorics, which is often used in mathematics to enumerate families of disjoint lattice paths. We refer the reader to [fromthebook, Chapter 29] for an accessible exposition of this theorem and some of its applications. For additional examples of this technique of pairing up and cancelling extraneous terms in matrix algebra, see [Zeilberger1985].
@@
Our exposition of the APC and $k$-APC algorithms works by interpreting certain determinants combinatorially, as generating functions for families of edge-disjoint walks. Previous work has also leveraged combinatorial interpretations of the determinant (as polynomials instead of formal power series however) to construct interesting arithmetic circuits [comb-det, det-hom]. In particular, the sign-reversing involution designed in [comb-det, Proof of Theorem 1] is very similar to the suffix swapping argument we use in the proof of Lemma 4.7.
The general approach of solving graph problems by testing whether certain enumerating polynomials are nonzero is now a common technique in graph algorithms: we refer the reader to [det-sieving] and the citations therein for previous examples of this paradigm in the literature.
@@
In the 4-Clique problem, we are given an undirected graph $G$ on $n$ nodes, and are tasked with determining if $G$ contains four vertices which are mutually adjacent. The best conditional lower bounds for APC come from observing that 4-Clique reduces to APC [ap-bounded-mincut, Section 4]. Using the current best algorithms for rectangular matrix multiplication [mmult-life, Table 1], the fastest known algorithm for 4-Clique takes $O(n^{3.251})$ time [EisenbrandGrandoni2004, Theorem 1]. Even if we conjecture this runtime time is optimal, the resulting lower bound for APC is a far cry from the $n^{4+o(1)}$ runtime we have for APC in dense graphs. Moreover, if $\omega=2$ this lower bound for APC weakens to $n^{3-o(1)}$.
@@
The fastest known deterministic verifier for APC runs in $O(n^{3.251}+n^{2.5}\sqrt{m})$ time [Trabelsi23]. It would be interesting to improve this runtime to $O(n^{3.251})$ to match the fastest known algorithm (and deterministic verifier runtime) for 4-Clique.
It would also be interesting to obtain faster randomized verifiers for APC. Currently, no randomized verifiers running faster than the deterministic verifier discussed above are known for APC. This is surprising, since the best lower bound for APC comes from the 4-Clique problem, and the 4-Clique problem admits a randomized verifier running in near-optimal $\tilde{O}(n^{2})$ time [MA-protocol, Section 4].
@@
In directed graphs, APVC and $k$-APVC reduce to APC and $k$-APVC respectively, so it might be easier to design faster algorithms for the former problems instead of the latter problems. For example, it is known that $k$-APVC can be solved in $\tilde{O}(k^{2}n^{\omega})$ time [k-APC, Theorem 5], which is faster than the $\tilde{O}((kn)^{\omega})$ runtime for $k$-APC if $\omega>2$.
In undirected graphs, although APC can be solved in $\tilde{O}(n^{2})$ time, APVC is not known to admit a near-quadratic time algorithm. In fact, the best known lower bounds for APC in directed graphs also hold for APVC in undirected graphs [vertex-conn-lower-bound].
So again, it might be easier to find faster algorithms for APVC in undirected graphs instead of tackling the general problem of APC in directed graphs. For example, APVC in undirected graphs can be solved in $\tilde{O}(m^{2})$ time [Trabelsi23, Section 3], which is faster than the $\tilde{O}(m^{\omega})$ runtime for APC if $\omega>2$, and admits a deterministic $O(n^{3.251})$ time verifier [Trabelsi23, Lemma 2.2], which is faster than the known $O(n^{3.521}+n^{2.5}\sqrt{m})$ time deterministic verifier for APC.